#11
Variable operating costs 45,000 Gross profit Fixed operating costs 20,000) Net operating income25,000 Interest Earnings before taxes 15,000) 10,000 4,000) 6,000 Taxes (40%) Net income Compute Surfsides degree of operating leverage (DOL), degree of financial leverage (DFL), and degree of total leverage (DTL). 12-11 Data Recovery Systems (DRS) has a degree of operating leverage (DOL) equal to 3.2x and a degree of total leverage (DTL) equal to 8x. DRS forecasts that this years sales will be $300,000 and that EBIT will be $180,000. If sales actually turn out to be $313,500, what will DRSs EBIT be? 12-12 The CFO of Jupiter Jibs (JJ) expects this years sales to be $2.5 million. EBIT is expected to be $1 million. The CFO knows that if sales actually turn out to be $2.3 million, JJs EBIT will be $880,000. What is JJs degree of operating leverage (DOL)? nTh crofTho chon shon has determined that the
Answer #1

# 11

DOL = 3.2

% change in sales = (New sales- original sales)/ original sales

= (313500-300000)/ 300000

=4.5%

DOL= % change in EBIT/ % change in sales

3.2 = % change in EBIT/ 4.5%

% change in EBIT= 3.2*4.5% = 14.4%

So new EBIT = Old EBIT*(100+% change in EBIT)

= 180000*(114.4%)

= 205920